Integrals Giving Inverse Trigonometric Functions


1. ∫(dx)/(x²+4x+5)


2. ∫(dt)/(t²-t+2)




1
Expert's answer
2021-04-13T15:43:59-0400

"1)x^2 + 4x +5 = (x+2)^2 +1"

So we have that

"\int \dfrac{dx}{x^2 +4x+5} = \int \dfrac{dx}{(x+2)^2 +1} = \arctan (x+2) + c"

where c is a constant

2) "t^2-t+2 = (t-\frac{1}{2})^2 + \frac{7}{4} = (t-\frac{1}{2})^2 + (\frac{\sqrt7}{2})^2"

So we have that

"\int \dfrac{dt}{t^2-t+2} = \int \dfrac{dt}{(t-\frac{1}{2})^2 + (\frac{\sqrt7}{2})^2} = \frac{1}{\frac{\sqrt7}{2}} \arctan(\frac{t- \frac{1}{2}}{\frac{\sqrt7}{2}}) + b" = "\frac{2}{\sqrt7} \arctan(\frac{2t-1}{\sqrt7}) + b"

where b is a constant